OpenStudy (supercalifragalisticexspeaalli):

who knows about Arithmetic series???????????? will medal & fan

OpenStudy (agent0smith):

A lot of people

OpenStudy (supercalifragalisticexspeaalli):

r u

OpenStudy (agent0smith):

Just ask the question already lol

OpenStudy (supercalifragalisticexspeaalli):

lol k

OpenStudy (supercalifragalisticexspeaalli):

Mike started a savings account by depositing $9. Each month, he deposits more money than the month before. At the end of 41 months, he has saved $9,389.00. How much more does he deposit each month? $11.00 $11.50 $12.00 $12.50

OpenStudy (supercalifragalisticexspeaalli):

?

OpenStudy (agent0smith):

hopefully you have seen this\[\large a_n = a_1 + (n-1)d\]a1 is first term, d is common difference and \[\large S_n = \frac{ n(a_1+a_n) }{ 2 }\]where Sn is the sum of the first n terms

OpenStudy (supercalifragalisticexspeaalli):

OH SO HOW WOULD WE DO THT

OpenStudy (agent0smith):

started a savings account by depositing $9 - so a1 = 9 end of 41 months - so n =41 he has saved $9,389.00 - so Sn = 9,389 How much more does he deposit each month? - we need to find d First, plug all those numbers into the second formula, the Sn one

OpenStudy (supercalifragalisticexspeaalli):

ok

OpenStudy (supercalifragalisticexspeaalli):

so how would that even look like

OpenStudy (agent0smith):

\[\large S_n = \frac{ n(a_1+a_n) }{ 2 }\] plug in a1 = 9 n =41 Sn = 9,389

OpenStudy (supercalifragalisticexspeaalli):